How much work is done when a 185 g tomato is lifted 10.6 m ?

The tomato is dropped. What is the velocity, v, of the tomato when it hits the ground? Assume 90.9 % of the work done in Part A is transferred to kinetic energy, E, by the time the tomato hits the ground.

Express your answer with the appropriate units.
Answer #1

A)

M = 185 g = 0.185 Kg

g = 9.81 m/s^2

h = 10.6 m

Use:

E = m*g*h

= 0.185 Kg * 9.81 m/s^2 * 10.6 m

= 19.2 J

Answer: 19.2 J

B)

KE = 90.9 % of 19.2 J

= 90.9*19.2/100

= 17.45 J

Use:

KE = 0.5*m*v^2

17.45 = 0.5*0.185*v^2

v^2 = 189 m/s

Answer: 189 m/s